OpenStudy (anonymous):

The line of symmetry of the parabola whose equation is y = ax2 - 4x + 3 is x = -2. What is the value of "a"?

OpenStudy (dumbcow):

line of symmetry is given by the formula: x=-b/2a -2 = 4/2a -4a = 4 a = -1